# Quantum Shortcut to Adiabaticity for State Preparation in a Finite-Sized Jaynes-Cummings Lattice

### v1

@@ III.2 Three sites one excitation
$H_{1}$ $=$ $ig_{n}\left[\left(a_{1}^{\dagger}+a_{3}^{\dagger}\right)\sigma_{2-}+a_{2}^{\dagger}\left(\sigma_{1-}+\sigma_{3-}\right)+h.c.\right]$ (12)
$+ig_{l}\left(a_{1}^{\dagger}\sigma_{3-}+a_{3}^{\dagger}\sigma_{1-}+h.c.\right)$
$+ig_{l}\left(a_{1}^{\dagger}\sigma_{1-}+2a_{2}^{\dagger}\sigma_{2-}+a_{3}^{\dagger}\sigma_{3-}+h.c.\right)$
which includes nonlocal couplings with strength $g_{n}$ or $g_{l}$ between different sites, such as the coupling between the cavity on sites $1$ and $3$ and the qubit on site $2$, as well as local couplings with strengths $g_{l}$ for sites $1$ and $3$, and strength $2g_{l}$ for site $2$.
We employ the same approach as in Sec. III.1 to simplify the Hamiltonian (12). As discussed earlier, when the initial state is $|v_{1}\rangle$, the ramping of the hopping terms induces transitions only to the state $|v_{4}\rangle$. Therefore, the coefficient $\alpha_{2}$ in $H_{1}$ can be varied with flexibility without affecting the evolution of the quantum state. By adjusting $\alpha_{2}$, we derive a simplified CD Hamiltonian in which the couplings between the qubit (cavity) on site $2$ and the cavities (qubits) on sites $1$ and $3$ are no longer required. The simplified CD Hamiltonian takes the form:
$H_{1}^{\prime}$ $=$ $ig^{\prime}\left(a_{1}^{\dagger}\sigma_{1-}+2a_{2}^{\dagger}\sigma_{2-}+a_{3}^{\dagger}\sigma_{3-}+h.c.\right)$ (13)
$+ig^{\prime}\left(a_{1}^{\dagger}\sigma_{3-}+a_{3}^{\dagger}\sigma_{1-}+h.c.\right),$
which includes nonlocal terms between sites $1$ and $3$ with coupling strength $g^{\prime}$. This Hamiltonian can be constructed for a small lattice of three sites by connecting the qubits and cavities on sites $1$ and $3$. However, this result shows that we cannot always simplify the CD Hamiltonian to only contain local couplings
@@ III.3 Two sites two excitations
In Fig. 4(a), we plot the eigenenergies vs the hopping rate $J$ at $g=1$ and $\Delta=0$. We denote the eigenstates as $|v_{i}\rangle$, with $i\in[1,8]$, corresponding to the eigenvalues given in (III.3). Here, $|v_{7}\rangle$ is the ground state, while $|v_{1}\rangle$ and $|v_{2}\rangle$ are degenerate states with $E_{1,2}=0$.
@@
Figure 4: (a) Eigenenergies $E_{n}$ vs hopping rate $J$. (b) Infidelity $1-F(t)$ vs $t/T$ for total evolution time $T=0.5\pi$. All solid lines: adiabatic evolution under the Hamiltonian $H_{\rm r}$; and all circles: adiabatic evolution with the CD Hamiltonian $H_{1}$ or $H_{1}^{\prime}$ applied. Blue: the initial state is the ground state $|v_{7}\rangle$ at $J=0$, and the CD Hamiltonian is $H_{1}$ for the full Hilbert space; red (green): the initial state is the ground state $|v_{7}\rangle$ (the first excited state $|v_{3}\rangle$) at $J=0$, and the CD Hamiltonian is $H_{1}^{\prime}$ for the subspace of $\{|v_{2}\rangle$, $|v_{3}\rangle$, and $|v_{4}\rangle\}$. Here, $g\equiv 1$, $\Delta\equiv 0$, $J(0)=0$, and $J_{f}=1$. The blue solid line is obscured completely by the green solid line.
@@
Ramping of hopping rate $J$. Assuming that the hopping rate is linearly ramped with $J=J_{f}t/T$, the target hopping rate $J_{f}=1$, $g\equiv 1$, and $\Delta\equiv 0$. The CD Hamiltonian for this process can be obtained numerically using Eq.(1). We simulate the dynamics of this system with the initial state being the ground state $|v_{7}\rangle$ for $J=0$ and with a total time $T=0.5\pi$. In Fig. 4(b), we plot the infidelity $1-F(t)$ of the quantum state at time $t$ as a function of $t/T$. The infidelity at the final time $T$ is above $0.4$ for the adiabatic evolution under the Hamiltonian $H_{\rm r}$, whereas it remains negligible when the CD Hamiltonian is applied.
@@
By examining the eigenstates of $H_{\rm r}$, we find that the eight eigenstates can be divided into two subsets. The first subset contains three states: ${|v_{2}\rangle,\,|v_{3}\rangle,\,|v_{4}\rangle}$, while the second subset contains the remaining five states, including the ground state $|v_{7}\rangle$. When varying the hopping rate $J$, the only nonzero transition matrix elements are between states within the same subset; there are no diabatic transitions between states from different subsets. When the initial state is from the first subset, the relevant diabatic transitions occur only within that subset. Thus, the CD Hamiltonian can be simplified by retaining only the counter-diabatic terms for the first subset in $H_{1}$, leading to the construction of a new CD Hamiltonian with these terms. Using this approach, we obtain a simplified CD Hamiltonian:
$H_{1}^{\prime}(t)=-i\frac{g}{2\chi_{2}^{2}}\frac{dJ}{dt}\left(A_{2}^{\dagger}C_{2}-C_{2}^{\dagger}A_{2}\right),$ (16)
@@
where $A_{2}=a_{1}^{2}-a_{2}^{2}$ and $C_{2}=a_{2}\sigma_{1-}-a_{1}\sigma_{2-}$. This Hamiltonian consists of four-operator terms, which can be formed through effective second-order couplings in the perturbative regime. While the engineering of such couplings has been previously studied, it’s important to note that the effective coupling strength may be significantly smaller compared to that of the first-order couplings $g$ and $J$.
@@
To demonstrate the effectiveness of Eq. (16), we conducted a numerical simulation of the adiabatic process using different initial states. In Fig. 4(b), when the initial state is $|v_{3}\rangle$ from the first subset, which corresponds to the first excited state as shown in Fig. 4(a), the infidelity remains below $10^{-10}$. This indicates that the diabatic transitions are completely cancelled by the CD Hamiltonian, preserving the system in the instantaneous eigenstate $|v_{3}\rangle$ throughout the evolution. However, when the initial state is $|v_{7}\rangle$ from the second subset, corresponding to the ground state of the system, the infidelities for the evolution with and without the CD terms are identical, suggesting that the CD Hamiltonian $H_{1}^{\prime}$ cannot cancel the diabatic transitions within the second subset. These numerical findings are consistent with our analysis of the eigenstates and the CD Hamiltonian.
@@ IV Error, Noise, and Measurement
In this section, we discuss the robustness of the CD driving against control errors and environmental noise, as well as the characterization of the CD approach through qubit measurements.
@@
Control errors. Adjusting the system parameters can introduce classical control errors, potentially impacting the performance of the CD approach. To investigate the effects of these errors, we conducted numerical simulations of the adiabatic process by introducing random fluctuations to the time-dependent coupling constants. Consider the linear ramping of the hopping rate $J(t)$ discussed in Sec. III.1, where the CD Hamiltonian only consists of time-dependent local qubit-cavity couplings. We introduced a fluctuation $\delta J(t)$ to the hopping rate, modifying it as $J(t)\rightarrow J(t)+\delta J(t)$, and a fluctuation $\delta g(t)$ to the qubit-cavity coupling [including the CD terms in (10)] with $g(t)\rightarrow g(t)+\delta g(t)$. Here, $\delta J(t)$ and $\delta g(t)$ are Gaussian random numbers with a standard deviation $\alpha J_{f}$ and $\alpha g$, respectively, with $\alpha$ being the ratio of the fluctuation magnitude to the coupling strength. We choose $\alpha$ within the range of $[0.01,0.15]$, i.e., the fluctuations are within $1\%$ $15\%$ of the coupling strengths. In Fig. 5(a), the fidelity $F(t)$ is plotted vs $t/T$ for a single sample of errors with $\alpha=0.05$. The fidelity exhibits small fluctuations over time due to the presence of errors. In Fig. 5(b), the fidelity $F(T)$ at time $T$, averaged over $100$ samples of errors, is plotted vs $\alpha$. The numerical result shows only a mild decrease in fidelity with the increase of $\alpha$, even for $\alpha\sim 15\%$, demonstrating the robustness of this process against control errors.
@@
Figure 5: (a) Fidelity $F(t)$ vs $t/T$ for one sample of errors with $\alpha=0.05$. (b) Fidelity $F(T)$ averaged over 100 samples vs $\alpha$. In (a) and (b), all blue circles: adiabatic evolution under the Hamiltonian $H_{\rm r}$; all red circles: adiabatic evolution with the CD Hamiltonian $H_{1}^{\prime}$ applied. The total evolution time $T=0.5\pi$. (c) Infidelity $1-F(T)$ vs $\gamma$ for $\kappa=5\times 10^{-5}$. (d) Infidelity $1-F(T)$ vs $\kappa$ for $\gamma=\frac{5}{\pi}\times 10^{-5}$. In (c) and (d), all solid lines: adiabatic evolution under the Hamiltonian $H_{\rm r}$; and all circles: adiabatic evolution with the CD Hamiltonian $H_{1}^{\prime}$ applied. Blue: $T=0.5\pi$; red: $T=3\pi$; and green: $T=5.5\pi$. All other parameters are the same as in Fig. 3.
@@
Decoherence. Environmental noise induces decoherence in the qubits and cavity modes, potentially impacting the fidelity of the prepared quantum states. To analyze the effects of decoherence in this system, we utilize a master equation approach. The master equation takes the form
@@
where $\mathcal{L}_{qj}=\frac{1}{2}(2\sigma_{j-}\rho\sigma_{j+}-\rho\sigma_{j+}\sigma_{j-}-\sigma_{j+}\sigma_{j-}\rho)$ is the Liouvillian operator for the qubit on site $j$ with damping rate $\gamma_{j}$, $\mathcal{L}_{aj}=\frac{1}{2}(2a_{j}\rho a_{j}^{\dagger}-\rho a_{j}^{\dagger}a_{j}-a_{j}^{\dagger}a_{j}\rho)$ is the Liouvillian operator for the cavity mode on site $j$ with damping rate $\kappa_{j}$, and $H_{\rm tot}$ is the total Hamiltonian of the system. Also, we choose $\gamma_{1,2}=\gamma$ and $\kappa_{1,2}=\kappa$ for simplicity of discussion. Assume that the dimensionless coupling strength $g=1$ in our discussion corresponds to $g=2\pi\times 100$ MHz in superconducting systems [55, 56]. The couplings in this system can be controlled by various methods, as demonstrated in recent experiments [57, 58, 59, 60, 61, 62, 63]. The measured decoherence time of $\sim 100{\rm\mu s}$ then corresponds to $\gamma=\frac{5}{\pi}\times 10^{-5}$ in dimensionless units. When the cavity frequency is $\omega_{0}=2\pi\times 5$ GHz and the quality factor is $Q=10^{6}$, the cavity decay rate is $\kappa=5\times 10^{-5}$ in dimensionless units.
@@
Figure 6: Operator average $\langle\sigma_{1x}\sigma_{2x}\rangle$ vs $t/T$ for $T=0.5\pi$. (a) One excitation with $g\equiv 1$, $\Delta\equiv 1$, and $J_{f}=2$ as in Fig. 3(a). (b) Two excitations with $g\equiv 1$, $\Delta\equiv 0$, and $J_{f}=1$ as in Fig. 4(a). Solid blue lines: adiabatic evolution under the Hamiltonian $H_{\rm r}$; blue circles: adiabatic evolution with the CD Hamiltonian $H_{1}^{\prime}$ applied; and solid red lines: instantaneous ground state average of Hamiltonian $H_{\rm r}$.
Measurement. The performance of the CD approach can be characterized by conducting measurement on the qubit operator $\sigma_{1x}\sigma_{2x}$. In the case of one excitation in a two-site lattice, the ground state at the initial parameters $g=1$, $J=0$, and $\Delta=1$ is $\frac{1}{2}(|1,-\rangle_{1}|0,g\rangle_{2}+|0,g\rangle_{1}|1,-\rangle_{2})$, where $|1,-\rangle$ is the lower-polariton state with one excitation, and the excitation occupies the sites in the form of a quantum superposition with the operator average $\langle\sigma_{1x}\sigma_{2x}\rangle=0.7236$. When the hopping rate becomes $J_{f}=2$ at time $T$, the operator average of the ground state decreases to $\langle\sigma_{1x}\sigma_{2x}\rangle=0.2764$, as shown in Fig. 6(a). It is noticeable that the operator average for the adiabatic evolution under $H_{\rm r}$ can significantly deviate from that of the ground state, while the operator average with the CD term applied remains the same as that of the ground state. Similarly, in the two-excitation case, the ground state at $g=1$, $J=0$, and $\Delta=0$ is $v_{7}=|1,-\rangle_{1}|1,-\rangle_{2}$, where each site is occupied by one excitation in the lower-polariton state with the operator average $\langle\sigma_{1x}\sigma_{2x}\rangle=0$. When the hopping rate becomes $J_{f}=1$, the operator average increases to $\langle\sigma_{1x}\sigma_{2x}\rangle=0.3659$. During the evolution, the operator average for the adiabatic process without the CD driving shows a large discrepancy with the operator average in the CD approach, as shown in Fig. 6(b). Hence, the effectiveness of the CD approach can be verified by measuring the operator average.

### v2

@@ III.2 Three sites one excitation
$H_{1}$ $=$ $ig_{n}\left[\left(a_{1}^{\dagger}+a_{3}^{\dagger}\right)\sigma_{2-}+a_{2}^{\dagger}\left(\sigma_{1-}+\sigma_{3-}\right)-h.c.\right]$ (12)
$+ig_{l}\left(a_{1}^{\dagger}\sigma_{3-}+a_{3}^{\dagger}\sigma_{1-}-h.c.\right)$
$+ig_{l}\left(a_{1}^{\dagger}\sigma_{1-}+2a_{2}^{\dagger}\sigma_{2-}+a_{3}^{\dagger}\sigma_{3-}-h.c.\right)$
which includes nonlocal couplings with strengths $g_{n}$ and $g_{l}$, such as the coupling between the cavities on sites $1$ and $3$ and the qubit on site $2$, as well as local couplings with strength $g_{l}$ for sites 1, 3131,\,31 , 3 and strength $2g_{l}$ for site $2$.
We employ the same approach as in Sec. III.1 to convert the CD Hamiltonian to a form that is easier to implement. As discussed above, for initial state $v_{1}$, ramping of the hopping terms only induces transition to the state $v_{4}$. Therefore, the coefficient $\alpha_{2}$ in $H_{1}$ can be varied with flexibility without affecting the dynamics of this system. By adjusting $\alpha_{2}$, we find
$H_{1}^{\prime}$ $=$ $ig^{\prime}\left(a_{1}^{\dagger}\sigma_{1-}+2a_{2}^{\dagger}\sigma_{2-}+a_{3}^{\dagger}\sigma_{3-}-h.c.\right)$ (13)
$+ig^{\prime}\left(a_{1}^{\dagger}\sigma_{3-}+a_{3}^{\dagger}\sigma_{1-}-h.c.\right),$
where the couplings between the qubit (cavity) on site $2$ and the cavities (qubits) on sites $1$ and $3$ are no longer required. This Hamiltonian still includes nonlocal terms between sites $1$ and $3$ with coupling strength $g^{\prime}$. This Hamiltonian can be implemented for a small lattice of three sites by coupling the qubits and cavities on sites $1$ and $3$. However, this result shows that we cannot always find a local form of the CD driving. In Fig. 3(c, d), we plot the infidelity of the prepared state under the adiabatic Hamiltonian $H_{\rm r}$ both without and with the CD driving $H_{1}^{\prime}$. The result is similar to that of the two-site lattice discussed in Sec. III.1.
@@ III.3 Two sites two excitations
In Fig. 4(a), we plot the eigenenergies vs the hopping rate $J$ at $g=1$ and $\Delta=0$. We denote the eigenstates as $v_{i}$ with $i\in[1,8]$, corresponding to the eigenvalues given in (15). Here, $v_{7}$ is the ground state, while $v_{1}$ and $v_{2}$ are degenerate states with $E_{1,2}=0$.
@@
Figure 4: (Color online) (a) Eigenenergies $E_{n}$ ($n\in[1,8]$) vs hopping rate $J$. (b) Infidelity $1-F(t)$ vs $t/T$ for a total evolution time of $T=0.5\pi$. Blue solid line (blue circles): linear ramping of hopping rate $J$ under the adiabatic Hamiltonian $H_{\rm r}$ only (with the CD driving $H_{1}$ applied) with the initial state being the ground state $v_{7}$ at $J=0$; and red dashed line (red squares): linear ramping of $J$ under $H_{\rm r}$ (with $H_{1}^{\prime}$) with the initial state being the first excited state $v_{3}$ at $J=0$. Here, $g\equiv 1$, $\Delta\equiv 0$, $J(0)=0$, $J_{f}=1$, and all parameters are in dimensionless units.
@@
Ramping of hopping rate $J$. Assuming that the hopping rate is linearly ramped with $J=J_{f}t/T$, the target hopping rate $J_{f}=1$, $g\equiv 1$, and $\Delta\equiv 0$. The CD driving $H_{1}$ can be obtained numerically using Eq. (1). We simulate the dynamics of this system with the initial state being the ground state $v_{7}$ at $J=0$ for a total evolution time of $T=0.5\pi$. In Fig. 4(b), we plot the infidelity $1-F(t)$ of the quantum state at time $t$ as a function of $t/T$. The infidelity at the final time $T$ exceeds $0.4$ under the adiabatic Hamiltonian $H_{\rm r}$, whereas it remains negligible when the CD driving is applied.
@@
By examining the eigenstates of $H_{\rm r}$, we find that the eight eigenstates can be divided into two subsets. The first subset contains three states $\{v_{2},\,v_{3},\,v_{4}\}$; while the second subset contains the remaining five states, including the ground state $v_{7}$. When varying the hopping rate $J$, the only nonzero transition matrix elements are between states within the same subset, and there is no diabatic transition between states in different subsets. Thus, for an initial state in the first subset, we consider the CD driving $H_{1}^{\prime}$ that only eliminates the diabatic transitions between eigenstates in this subset. As detailed in Appendix B, we derive
$H_{1}^{\prime}(t)=i\frac{g}{2\left(J^{2}+2g^{2}\right)}\frac{dJ}{dt}\left(A_{2}^{\dagger}C_{2}-C_{2}^{\dagger}A_{2}\right),$ (16)
@@
where $A_{2}=a_{1}^{2}-a_{2}^{2}$ and $C_{2}=a_{2}\sigma_{1-}-a_{1}\sigma_{2-}$. This Hamiltonian consists of four-operator terms, which can be formed through effective second-order couplings in the perturbative regime.
@@
To demonstrate the effectiveness of Eq. (16), we conducted a numerical simulation of the adiabatic process on the initial state $v_{3}$ from the first subset, which corresponds to the first excited state in Fig. 4(a). As shown in Fig. 4(b), up to a small numerical error below $10^{-10}$, the infidelity remains zero throughout the evolution, indicating that the diabatic transitions are completely cancelled by the CD driving $H_{1}^{\prime}$. In contrast, if the initial state is the ground state $v_{7}$ from the second subset, the infidelities both without and with the CD driving $H_{1}^{\prime}$ will be the same [see the blue solid line in Fig. 4(b)]. These numerical findings clearly show that the CD driving $H_{1}^{\prime}$ is only effective for states in the first subset, but not for states in the second subset, which is consistent with our analysis above.
@@ IV Error, Noise, and Measurement
In this section, we discuss the effects of control errors and environmental noise on the adiabatic evolution under the CD driving, as well as the characterization of the CD approach through qubit measurements.
@@
Control errors. Adjusting the system parameters can introduce classical control errors, potentially impacting the performance of the CD approach. To investigate the effects of these errors, we conducted numerical simulations of the adiabatic process by introducing random fluctuations to the time-dependent coupling constants. Consider the linear ramping of the hopping rate $J(t)$ discussed in Sec. III.1, where the CD driving only consists of time-dependent local qubit-cavity couplings. We introduce a fluctuation $\delta J(t)$ to the hopping rate with $J(t)\rightarrow J(t)+\delta J(t)$, and a fluctuation $\delta g(t)$ to the qubit-cavity coupling [including the CD terms in (10)] with $g(t)\rightarrow g(t)+\delta g(t)$. Here, $\delta J(t)$ and $\delta g(t)$ are Gaussian random numbers with a standard deviation $\alpha J_{f}$ and $\alpha g$, respectively, where $\alpha$ is the ratio of the fluctuation amplitude to the coupling strength. We choose $\alpha$ to be below $0.15$, i.e., the fluctuations are below $15\%$ of the coupling strengths. In superconducting quantum devices, the couplings can be controlled by various methods, as demonstrated in recent experiments [66, 67, 68, 69, 70, 71, 72], and the ratio $\alpha$ can be well below $5\%$ [73, 74]. In Fig. 5(a), the fidelity $F(t)$ is plotted vs $t/T$ for a single sample of errors with $\alpha=0.05$ and $T=0.5\pi$. The fidelity exhibits small fluctuations over the time $t$ due to the presence of the errors. In Fig. 5(b), the fidelity $F(T)$ averaged over $100$ samples of errors is plotted vs the ratio $\alpha$. As $\alpha$ increases to $15\%$, the fidelity only shows a mild decrease, demonstrating that control errors will not strongly affect the system dynamics in practical systems.
@@
Figure 5: (Color online) (a) Fidelity $F(t)$ vs $t/T$ for one sample of errors with $\alpha=0.05$. (b) Fidelity $F(T)$ vs $\alpha$ averaged over 100 samples of errors. In (a) and (b), blue circles (red triangles): under the adiabatic Hamiltonian $H_{\rm r}$ only (with the CD driving $H_{1}^{\prime}$ applied); and the total evolution time is $T=0.5\pi$. (c) Infidelity $1-F(T)$ vs $T$ for $\gamma=\frac{5}{\pi}\times 10^{-5}$. Blue solid line (blue circles): for $\kappa=5\times 10^{-4}$ under the adiabatic Hamiltonian $H_{\rm r}$ only (with the CD driving $H_{1}^{\prime}$ applied); red dashed line (red squares): for $\kappa=10^{-4}$ under $H_{\rm r}$ (with $H_{1}^{\prime}$); and green dotted line (green triangles): for $\kappa=5\times 10^{-5}$ under $H_{\rm r}$ (with $H_{1}^{\prime}$). (d) Infidelity $1-F(T)$ vs $\kappa$ for $\gamma=\frac{5}{\pi}\times 10^{-5}$. Blue solid line (blue circles): for $T=0.5\pi$ under the adiabatic Hamiltonian $H_{\rm r}$ only (with the CD driving $H_{1}^{\prime}$ applied); red dashed line (red squares): for $T=3\pi$ under $H_{\rm r}$ (with $H_{1}^{\prime}$); and green dotted line (green triangles): for $T=5.5\pi$ under $H_{\rm r}$ (with $H_{1}^{\prime}$). All plots are for linear ramping of hopping rate $J$ with parameters given in Fig. 3 in dimensionless units.
@@
Decoherence. Environmental noise induces decoherence in the qubits and the cavity modes, potentially impacting the fidelity of the prepared quantum states. To analyze the effects of decoherence, we utilize a master equation approach with
@@
where $\mathcal{L}_{qj}=\frac{1}{2}(2\sigma_{j-}\rho\sigma_{j+}-\rho\sigma_{j+}\sigma_{j-}-\sigma_{j+}\sigma_{j-}\rho)$ is the Liouvillian operator for the qubit on site $j$ with damping rate $\gamma_{j}$, $\mathcal{L}_{aj}=\frac{1}{2}(2a_{j}\rho a_{j}^{\dagger}-\rho a_{j}^{\dagger}a_{j}-a_{j}^{\dagger}a_{j}\rho)$ is the Liouvillian operator for the cavity mode on site $j$ with damping rate $\kappa_{j}$, and $H_{\rm tot}$ is the total Hamiltonian of the system. We choose $\gamma_{1,2}=\gamma$ and $\kappa_{1,2}=\kappa$ for simplicity of discussion. We also assume that the dimensionless coupling strength $g=1$ in our discussion corresponds to $g=2\pi\times 100$ MHz in superconducting systems [62, 63]. The measured decoherence time of $\sim 100{\rm\mu s}$ then corresponds to $\gamma=\frac{5}{\pi}\times 10^{-5}$ in dimensionless units. When the cavity frequency is $\omega_{0}=2\pi\times 5$ GHz and the quality factor is $Q=10^{6}$, the cavity decay rate is $\kappa=5\times 10^{-5}$ in dimensionless units.
@@
Figure 6: (Color online) Operator average $\langle\sigma_{1x}\sigma_{2x}\rangle$ vs $t/T$ for $T=0.5\pi$. (a) Two sites one excitation with linear ramping of $J$ and parameters given in Fig. 3 in dimensionless units. (b) Two sites two excitations with linear ramping of $J$ and parameters given in Fig. 4 in dimensionless units. In (a) and (b), blue solid line (blue circles): under the adiabatic Hamiltonian $H_{\rm r}$ only (with the CD driving $H_{1}^{\prime}$ applied); and red dashed line: instantaneous ground state average of $H_{\rm r}$.
Measurement. The performance of the CD approach can be characterized by conducting measurement on the qubit operator $\sigma_{1x}\sigma_{2x}$. In the case of one excitation in a two-site lattice, the ground state at the initial parameters $g=1$, $J=0$, and $\Delta=1$ is $\frac{1}{2}(|1,-\rangle_{1}|0,g\rangle_{2}+|0,g\rangle_{1}|1,-\rangle_{2})$, where $|1,-\rangle$ is the lower-polariton state with one excitation, and the excitation occupies the sites in the form of a quantum superposition with the operator average $\langle\sigma_{1x}\sigma_{2x}\rangle=0.7236$. When the hopping rate becomes $J_{f}=2$ at time $T$, the operator average of the ground state decreases to $\langle\sigma_{1x}\sigma_{2x}\rangle=0.2764$, as shown in Fig. 6(a). It is noticeable that the operator average for the adiabatic evolution under $H_{\rm r}$ can significantly deviate from that of the ground state, while the operator average with the CD driving applied remains the same as that of the ground state. Similarly, in the two-excitation case, the ground state at $g=1$, $J=0$, and $\Delta=0$ is $v_{7}=|1,-\rangle_{1}|1,-\rangle_{2}$, where each site is occupied by one excitation in the lower-polariton state with the operator average $\langle\sigma_{1x}\sigma_{2x}\rangle=0$. When the hopping rate becomes $J_{f}=1$, the operator average increases to $\langle\sigma_{1x}\sigma_{2x}\rangle=0.3659$. During the evolution, the operator average for the adiabatic process without the CD driving shows a large discrepancy with the operator average in the CD approach, as shown in Fig. 6(b). Hence, the effectiveness of the CD approach can be verified by measuring the operator average.